*My brother is 12 years old than my sister. In 3 years, my sister will be 25 years old. How old is my brother?

* A coat costs 165.000 more than a jacket. If the jacketīs price is 30% of the coatīs price, what is the price of the coat?

Answer by stanbon(75887)

My brother is 12 years older than my sister. In 3 years, my sister will be 25 years old. How old is my brother?
-----
b = s + 12
s + 3 = 25
----
s = 25-3 = 22 (sister's age now)
b = s + 12 = 22+12 = 34 yrs (brother's age now)
-------------------------
* A coat costs 165.000 more than a jacket. If the jacketīs price is 30% of the coatīs price, what is the price of the coat?
c = j + 165
j = 0.3*c
-------
Ans: c = 0.3c + 165
0.7c = 165
coat cost = 165/0.7 = $235.71
-------------------
Cheers,
Stan H.